OpenStudy (anonymous):

Over the first 8 games of the 2009 season, Brett Favre threw three interceptions. If he were to continue that pace, how many interceptions would he throw over the entire 16 game season?

OpenStudy (texaschic101):

3/8 = x/16 (3 interceptions for 8 games = x interceptions for 16 games) cross multiply 8x = 48 x = 48/8 x = 6 6 interceptions over the entire 16 game season